# The $\ell$-test: leveraging sparsity in the Gaussian linear model for improved inference

### v1

@@ C.3.1 Proof of continuity and increasing properties of $f_{\mathbf{S}_{j}}$
$\implies$ $-\bm{X}_{j}^{T}\bm{P}_{\mathcal{A}(a)}\bm{X}_{j}\frac{\partial}{\partial a}\hat{\beta}_{j}(a)+\|\bm{X}_{j}\|^{2}\frac{\partial}{\partial a}\hat{\beta}^{\lambda}_{j}(a)=\|(\bm{I}-\bm{P}_{-j})(\bm{X}_{j})\|^{2}$
$\implies$ $\frac{\partial}{\partial a}\hat{\beta}^{\lambda}_{j}(a)=\frac{\|(\bm{I}-\bm{P}_{-j})(\bm{X}_{j})\|^{2}}{\|\bm{X}_{j}\|^{2}-\|\bm{P}_{\mathcal{A}(a)}\bm{X}_{j}\|^{2}}=\frac{\|(\bm{I}-\bm{P}_{-j})(\bm{X}_{j})\|^{2}}{\|(\bm{I}-\bm{P}_{\mathcal{A}(a)})(\bm{X}_{j})\|^{2}},$
which is positive, whenever $\hat{\beta}^{\lambda}_{j}(a)\neq 0$. Hence, for an $a\in\mathbb{R}$, either $\hat{\beta}^{\lambda}_{j}(a)=0$ or $\frac{\partial\hat{\beta}_{j}(a)}{\partial a}>0$, showing that $\hat{\beta}^{\lambda}_{j}$ is locally increasing around $a$ in the latter case. Now define,
@@
and note that, $\hat{\beta}_{j}^{\lambda}=f_{\bm{S}^{(j)}}(u_{1})$. Furthermore, note that $f_{\bm{S}^{(j)}}$ is a functional of the sufficient statistic, $\bm{S}^{(j)}$, and using the arguments above, is a continuous, piece-wise linear function that is increasing in the region, $\{u:f_{\bm{S}^{(j)}}(u)\neq 0\}$. This establishes all the claims about $f_{\bm{S}^{(j)}}$ in the statement of Theorem 2.1 except for (2.4), which we turn to next.
@@
C.3.2 Proof of Equation (2.4) in Theorem 2.1
@@
Theorem C.1.
For data $(\bm{y},\bm{X})$, with $\bm{X}$ full column-rank and $\lambda>0$, define $\bm{\hat{\beta}}_{*}^{\lambda}(b)$ to be $b$ at the $j^{\mathrm{th}}$ coordinate and $\bm{\hat{\beta}}_{-j}^{\lambda}(b)$ on the rest, where, $\bm{\hat{\beta}}_{-j}^{\lambda}(b)$ is defined as in Equation (2.3). Also let
$f_{\lambda}(\bm{y};\bm{\beta}):=\frac{1}{2n}\|\bm{y}-\bm{X}\bm{\beta}\|^{2}+\lambda\|\bm{\beta}\|_{1},$ (C.8)
@@
Thus, if one runs a blockwise coordinate descent with blocks $\{j\}$ and $[1:p]\setminus\{j\}$ starting at $\hat{\bm{\beta}}_{*}^{\lambda}(b)$, we see that the iterates will be constant at $\hat{\bm{\beta}}_{*}^{\lambda}(b)$, thereby implying that this is a limit point of the iterates. One can now invoke (Tseng, 2001, Proposition 5.1) (the conditions for applying this proposition follow directly as $f_{\lambda}$ can be separated into the squared error loss and the non-differentiable $\ell_{1}$-penalty) to conclude that,
@@
Hence, Theorem C.1 now implies that $\hat{\beta}_{j}^{\lambda}=b$ if and only if $0\in\partial_{\beta_{j}}f_{\lambda}(\bm{y};\bm{\beta})\lvert_{\bm{\beta}=\hat{\bm{\beta}}_{*}^{\lambda}(b)}$. We will now prove item 2 of Theorem 2.1 by evaluating this sub-gradient.
Proof of item 2 of Theorem 2.1.
@@
From the calculations in Section C.2,
@@
Setting $\bm{\beta}=\hat{\bm{\beta}}^{*}(b)$ for a $b$, the above equation along with Theorem C.1 establishes that for $b\neq 0$, $\hat{\beta}_{j}^{\lambda}=b$ if and only if,
@@
C.4 Proof of the fact that $T_{j}$ is independent of the sufficient statistic, $\mathbf{S}_{j}$ under $H_{j}:\beta_{j}=0$
@@
In Section C.2, we showed that,

### v2

@@ D.3.1 Proof of continuity and increasing properties of $f_{\mathbf{S}_{j}}$
$\implies$ $-\bm{X}_{j}^{T}\bm{P}_{-j}\bm{P}_{\mathcal{A}(a)}\bm{P}_{-j}\bm{X}_{j}\frac{\partial}{\partial a}\hat{\beta}_{j}(a)+\|\bm{X}_{j}\|^{2}\frac{\partial}{\partial a}\hat{\beta}^{\lambda}_{j}(a)=\|(\bm{I}-\bm{P}_{-j})(\bm{X}_{j})\|^{2}$
$\implies$ $-\bm{X}_{j}^{T}\bm{P}_{-j}\bm{P}_{\mathcal{A}(a)}\bm{P}_{\mathcal{A}(a)}\bm{P}_{-j}\bm{X}_{j}\frac{\partial}{\partial a}\hat{\beta}_{j}(a)+\|\bm{X}_{j}\|^{2}\frac{\partial}{\partial a}\hat{\beta}^{\lambda}_{j}(a)=\|(\bm{I}-\bm{P}_{-j})(\bm{X}_{j})\|^{2}$
$\implies$ $-\|\bm{P}_{\mathcal{A}(a)}\bm{P}_{-j}\bm{X}_{j}\|^{2}\frac{\partial}{\partial a}\hat{\beta}_{j}(a)+\|\bm{X}_{j}\|^{2}\frac{\partial}{\partial a}\hat{\beta}^{\lambda}_{j}(a)=\|(\bm{I}-\bm{P}_{-j})(\bm{X}_{j})\|^{2}$
$\implies$ $\frac{\partial}{\partial a}\hat{\beta}^{\lambda}_{j}(a)=\frac{\|(\bm{I}-\bm{P}_{-j})(\bm{X}_{j})\|^{2}}{\|\bm{X}_{j}\|^{2}-\|\bm{P}_{\mathcal{A}(a)}\bm{P}_{-j}\bm{X}_{j}\|^{2}},$
which is positive (as $\|\bm{P}_{\mathcal{A}(a)}\bm{P}_{-j}\bm{X}_{j}\|^{2}\leq\|\bm{P}_{-j}\bm{X}_{j}\|^{2}\leq\|\bm{X}_{j}\|^{2}$), whenever $\hat{\beta}^{\lambda}_{j}(a)\neq 0$. Hence, for an $a\in\mathbb{R}$, either $\hat{\beta}^{\lambda}_{j}(a)=0$ or $\frac{\partial\hat{\beta}_{j}(a)}{\partial a}>0$, showing that $\hat{\beta}^{\lambda}_{j}$ is locally increasing around $a$ in the latter case. Now define,
@@
and note that, $\hat{\beta}_{j}^{\lambda}=f_{\bm{S}^{(j)}}(u_{1})$. Furthermore, note that $f_{\bm{S}^{(j)}}$ is a functional of the sufficient statistic, $\bm{S}^{(j)}$, and using the arguments above, is a continuous, piece-wise linear function that is increasing in the region, $\{u:f_{\bm{S}^{(j)}}(u)\neq 0\}$. This establishes all the claims about $f_{\bm{S}^{(j)}}$ in the statement of Theorem 1 except for (2.4), which we turn to next.
@@
D.3.2 Proof of Equation (2.4) in Theorem 1
@@
Theorem 5.
For data $(\bm{y},\bm{X})$, with $\bm{X}$ full column-rank and $\lambda>0$, define $\hat{\bm{\beta}}_{*}^{\lambda}(b)$ to be $b$ at the $j^{\mathrm{th}}$ coordinate and $\hat{\bm{\beta}}_{-j}^{\lambda}(b)$ on the rest, where, $\hat{\bm{\beta}}_{-j}^{\lambda}(b)$ is defined as in Equation (2.3). Also let
$f_{\lambda}(\bm{y};\bm{\beta}):=\frac{1}{2n}\|\bm{y}-\bm{X}\bm{\beta}\|^{2}+\lambda\|\bm{\beta}\|_{1},$ (D.8)
@@
Thus, if one runs a blockwise coordinate descent with blocks $\{j\}$ and $1:p]\setminus\{j\}$ starting at $\hat{\bm{\beta}}_{*}^{\lambda}(b)$, we see that the iterates will be constant at $\hat{\bm{\beta}}_{*}^{\lambda}(b)$, thereby implying that this is a limit point of the iterates. One can now invoke tseng2001 (the conditions for applying this proposition follow directly as $f_{\lambda}$ can be separated into the squared error loss and the non-differentiable $\ell_{1}$-penalty) to conclude that,
@@
Hence, Theorem 5 now implies that $\hat{\beta}_{j}^{\lambda}=b$ if and only if $0\in\partial_{\beta_{j}}f_{\lambda}(\bm{y};\bm{\beta})\lvert_{\bm{\beta}=\hat{\bm{\beta}}_{*}^{\lambda}(b)}$. We will now prove item 2 of Theorem 1 by evaluating this sub-gradient.
Proof of item 2 of Theorem 1.
@@
From the calculations in Section D.2,
@@
Setting $\bm{\beta}=\hat{\bm{\beta}}^{*}(b)$ for a $b$, the above equation along with Theorem 5 establishes that for $b\neq 0$, $\hat{\beta}_{j}^{\lambda}=b$ if and only if,
@@
D.4 Proof of the fact that $T_{j}$ is independent of the sufficient statistic, $\mathbf{S}_{j}$ under $H_{j}:\beta_{j}=0$
@@
In Section D.2, we showed that,